OpenStudy (anonymous):

Find a third-degree polynomial function such that f(0) = -12 and whose zeros are 1, 2, and 3. Using complete sentences, explain how you found it

OpenStudy (aravindg):

well if 1,2, 3 are the roots (x-1)(x-2)(x-3) will be the third degree polynomial

OpenStudy (aravindg):

but an additional condition given is that f(0)=-12

OpenStudy (aravindg):

so we will have to make a slight change in our above polynomial ..any ideas?

OpenStudy (aravindg):

note for (x-1)(x-2)(x-3) f(0)=-6

OpenStudy (anonymous):

wait i dont understand

OpenStudy (aravindg):

what do you not understand?

OpenStudy (anonymous):

why did you do this (x-1)(x-2)(x-3) f(0)=-6 ?

OpenStudy (aravindg):

do you understand how i got (x-1)(x-2)(x-3) ?

OpenStudy (anonymous):

because the domains are 1 2 3?

OpenStudy (aravindg):

because zeroes of function are 1 ,2,3

OpenStudy (anonymous):

ok i got that part.

OpenStudy (anonymous):

now wat?

OpenStudy (aravindg):

now we could have reported this answer but the question insists on one more condition ie f(0)=-12

OpenStudy (aravindg):

we currently have f(X)=(x-1)(x-2)(x-3) f(0) =?

OpenStudy (anonymous):

-12?

OpenStudy (aravindg):

dont be lazy put x=0 and tell me what you get !

OpenStudy (anonymous):

so it would be 0-1=-1? 0-2=-2? 0-3=-3?

OpenStudy (aravindg):

try again ?

OpenStudy (anonymous):

omg this is so hard! do i plug in 1 for f(0)? like f(1)?

OpenStudy (aravindg):

plug in 0 for f(0)

OpenStudy (anonymous):

i dont understand! if i plug in 0 it would be f(0)

OpenStudy (anonymous):

suppose f(X)=ax^3+bx^2+cx+d so f(0)=-12 --->put x=0 hten -12=d so d=-12 f(1)=0 as 1 is root a+b+c+d=0 now f(2)=0 8a+4b+2c+d=0 f(3)=0 27a+9b+3c+d=0 solve the system of equation

OpenStudy (anonymous):

i give up :/